0.400 L of HBr were neutralized with 0. 2 L of a 6 M KOH solution. What was the molarity of the acid neutralized?
Answer #1

The balanced equation is

KOH + HBr ------> KBr + H2O

Number of moles of KOH = molarity * volume of solution in L

Number of moles of KOH = 6 * 0.2 = 1.2 mole

From the balanced equation we can say that

1 mole of KOH requires 1 mole of HBr so

1.2 mole of KOH will require

= 1.2 mole of KOH *(1 mole of HBr / 1 mole of KOH)

= 1.2 mole of HBr

Molarity of HBr = number of moles of HBr / volume of solution in L

Molarity of HBr = 1.2 / 0.400 = 3 M

Therefore, the molarity of acid = 3 M